James Allen Morrow, Bibliography

[1] A survey of some results on complex Kähler manifolds, Global Analysis (1969), 315– 324, University of Tokyo Press.

[2] The tangent bundle of the long line, Proc. Amer. Math. Soc. 23 (1969), 458.

[3] The denseness of complete Riemannian metrics, J. Differential Geometry 4 (1970), 225–226.

[4] Complex Manifolds, vii + 192 pp., Holt,Rinehart(1971).

[5] Nonsingular deformations of singular spaces with normal crossings, Proceedings of Liverpool Singularities Symposium II, Springer Lecture Notes in Mathematics 209, (1971), 234–243.

[6] Compactifications of C^{2}, Bull. Amer. Math. Soc. 78 (1972), 813–816.

[7] Minimal normal compactifications of C^{2}, Rice University Studies (Conference on Complex Analysis, 1972) 59 (1), (1973), 97–112.

[8] Non-singular deformations of spaces with normal crossings, I, J. Differential Geometry 9 (1974), 195–218.

[9] Some theorems of algebraicity for complex spaces, J. Math. Soc. Japan 27 (1975), 167–183 (co-author Hugo Rossi).

[10] Compactifying C^{n}, Several Complex Variables, Williamstown, 1975, Proc. of Symposia in Pure Math. 30 (1977), part I, 241–246 (co-author Larry Brenton).

[11] Compactifications of C^{n}, Trans. Amer. Math. Soc. 246 (1976), 139–153 (co-author Larry Brenton).

[12] Submanifolds of P^{n} with splitting normal bundle sequence are linear, Math. Ann. 234 (1978), 253–261 (co-author Hugo Rossi).

[13] Some general results on equivalence of embeddings, Recent Developments in Several Complex Variables, (1981) Princeton University Press (co-author Hugo Rossi).

[14] Canonical embeddings, Trans. Amer. Math. Soc. 261 (1981), 547-565 (co-author Hugo Rossi).

[15] Pluri-canonical embeddings, Compositio Math. 44 (1981), 219-239 (co-author Hugo Rossi).

[16] Curves in a ruled surface, (co-author Hugo Rossi).

[17] Compositions of slit mappings, preprint (co-author Don Marshall).

[18] Calculating the Resistors in a Network, Proceedings of the Annual International Conference of the IEEE Engineering in Medicine and Biology Society 11, part 2 (1989), p 451-452 (co-author Ed Curtis and Thad Edens).

[19] Determining the Resistors is a Network, Siam J. Appl. Math 50(1990), 918-930 (coauthor Ed Curtis).

[20] The Dirichlet to Neumann Map for a resistor network, Siam J. Appl. Math 51(1991), 1011-1029 (co-author Ed Curtis).

[21] Finding the Conductors in a Circular Network from Boundary Measurements, Mathematical Modelling and Numerical Analysis 28 (1994), 781-813 (co-authors Ed Curtis and Edith Mooers).

[22] Circular Planar Graphs and Resistor Networks, preprint (co-authors Edward B. Curtis and David Ingerman).